# Redshift Dipoles from Non-Geodesic Observer Congruences in Covariant Cosmology

### 21 Mar 2026 v1

@@ 4 Non-geodesic observer congruence and the dipole modulation
$\frac{1}{1+z}\frac{d(1+z)}{d\lambda}=-EH+E\,\mathcal{A}_{\alpha}e^{\alpha}.$ (4.4)
At this stage we keep the FLRW background relation for the scale factor evolution 3.13, so that the modification arises entirely from the observer-congruence dependence of the redshift propagation law. We obtain
$\frac{d}{d\lambda}\ln\!\big[(1+z)a\big]=E\,\mathcal{A}_{\alpha}e^{\alpha}.$ (4.5)
Therefore, the standard relation $(1+z)a=a_{0}$ is replaced by
@@
$(1+z_{obs})=\frac{a_{0}}{a}\exp\left(\int_{\lambda_{0}}^{\lambda_{s}}E\,\mathcal{A}_{\alpha}e^{\alpha}\,d\lambda\right).$ (4.6)
@@
For weak non-geodesic effects, this becomes
@@
$1+z_{obs}\simeq\frac{a_{0}}{a}\left[1+\int_{\lambda_{0}}^{\lambda_{s}}E\,\mathcal{A}_{\alpha}e^{\alpha}\,d\lambda\right].$ (4.7)
@@
Since $\mathcal{A}_{\alpha}e^{\alpha}$ is the projection of a spatial vector along the line of sight, it has the angular structure of a dipole. This shows that non-geodesicity of the observer congruence naturally induces a dipolar modulation in the redshift itself. The non-geodesic dipole arises from a line-of-sight integral of the observer acceleration field and therefore depends on the photon trajectory and redshift. This contrasts with the standard kinematic dipole, which is generated by a constant boost and is independent of distance. As a result, this dipole cannot, in general, be removed by a single global Lorentz transformation, as it arises from an integrated, position-dependent acceleration field rather than a constant boost.
5 Order–of–Magnitude Estimate of the Non–Geodesic Dipole
The dipolar contribution associated with a non–geodesic observer congruence arises from the term $\mathcal{A}_{a}e^{a}$ in the covariant propagation equation of the redshift,
$\frac{d}{d\lambda}\ln(1+z)=-E\left(\frac{1}{3}\theta-\mathcal{A}_{a}e^{a}\right).$ (5.1)
A simple estimate can be obtained by isolating the non-geodesic contribution. We focus on the dipolar term alone and define
$\mathcal{A}_{\parallel}\equiv\mathcal{A}_{a}e^{a}.$ (5.2)
The shear term has already been neglected in the previous section, so that only the isotropic expansion term and the dipolar non–geodesic term remain. To estimate the amplitude of the anisotropic modulation, it is convenient to separate the isotropic monopole contribution as
$\ln(1+z)\simeq\ln(1+\bar{z})+\delta\ln(1+z)_{\rm ng}.$ (5.3)
where $\bar{z}$ is the associated monopolar redshift, whose evolution is
$\frac{d}{d\lambda}\ln(1+\bar{z})=-EH.$ (5.4)
The dipolar residual is therefore governed by
@@ 5 Order–of–Magnitude Estimate of the Non–Geodesic Dipole
$\left.\frac{d}{d\lambda}\,\delta\ln(1+z)\right|_{\rm ng}\simeq E\,\mathcal{A}_{\parallel},\qquad\mathcal{A}_{\parallel}\equiv\mathcal{A}_{a}e^{a}.$ (5.5)
@@
$k^{a}=\frac{dx^{a}}{d\lambda},$ (5.6)
and therefore, in the local rest frame of the observer congruence,
$E=-u_{a}k^{a}=\frac{dt}{d\lambda},$ (5.7)
up to the usual choice of units and sign convention. [^1] More precisely, in the local rest frame of the observer congruence one has $E=c\,dt/d\lambda$ if $c$ is kept explicit. Using
$\frac{d}{d\lambda}=\frac{dt}{d\lambda}\frac{d}{dt}=E\frac{d}{dt},$ (5.8)
Eq. (5.5) becomes
$\left.\frac{d}{dt}\delta\ln(1+z)\right|_{\rm ng}\simeq\mathcal{A}_{\parallel},$ (5.9)
in units with $c=1$. Restoring dimensions, the relevant rate is $\mathcal{A}_{\parallel}/c$, so that
$\left.\frac{d}{dt}\delta\ln(1+z)\right|_{\rm ng}\sim\frac{\mathcal{A}_{\parallel}}{c}.$ (5.10)
@@
$\delta\ln(1+z_{obs})_{\rm ng}\sim\int\frac{\mathcal{A}_{\parallel}}{c}\,dt.$ (5.11)
If $\mathcal{A}_{\parallel}$ varies slowly along the line of sight over the redshift range of interest, this reduces to
$\delta\ln(1+z_{obs})_{\rm ng}\approx\frac{\mathcal{A}_{\parallel}}{c}\,\Delta t\approx\frac{\mathcal{A}_{\parallel}L}{c^{2}},$ (5.12)
@@
$L\simeq\frac{c}{H_{0}}\,z.$ (5.13)
Substituting Eq. (5.13) into Eq. (5.12) yields
$\delta\ln(1+z_{obs})_{\rm ng}\sim\frac{\mathcal{A}_{\parallel}}{cH_{0}}\,z.$ (5.14)
For $z\ll 1$, one has $\delta\ln(1+z)\simeq\delta z$ to leading order, so the same estimate can be written directly as
$\delta z_{\rm ng}\sim\frac{\mathcal{A}_{\parallel}}{cH_{0}}\,z.$ (5.15)
Inverting this relation gives an estimate of the effective acceleration required to generate a dipole of amplitude $\delta z_{\rm ng}$ at redshift $z$:
$\mathcal{A}_{\parallel}\sim\frac{cH_{0}}{z}\,\delta z_{\rm ng}.$ (5.16)
@@
$cH_{0}\simeq 6.8\times 10^{-10}\ {\rm m\,s^{-2}}.$ (5.17)
At $z\simeq 0.1$, representative of the upper end of the shallow redshift regime relevant for CosmicFlows analyses (see [14]), Eq. (5.16) becomes
$\mathcal{A}_{\parallel}\simeq 6.8\times 10^{-9}\,\delta z_{\rm ng}\quad{\rm m\,s^{-2}}.$ (5.18)
Therefore, a dipole amplitude of $\delta z_{\rm ng}\sim 10^{-3}$, $\delta z_{\rm ng}\sim 10^{-4}$ and $\delta z_{\rm ng}\sim 10^{-5}$ requires, respectively
$\mathcal{A}_{\parallel}\sim 7\times 10^{-12}\ {\rm m\,s^{-2}},$ (5.19)
$\mathcal{A}_{\parallel}\sim 7\times 10^{-13}\ {\rm m\,s^{-2}},$ (5.20)
$\mathcal{A}_{\parallel}\sim 7\times 10^{-14}\ {\rm m\,s^{-2}}.$ (5.21)
For comparison, a characteristic cosmological acceleration associated with the buildup of a peculiar velocity $v\sim 300\,{\rm km\,s^{-1}}$ over a Hubble time is
$\mathcal{A}_{\rm H}\sim vH_{0}\sim 7\times 10^{-13}\ {\rm m\,s^{-2}},$ (5.22)
which lies in the range required to generate a dipole of order $\delta z\sim 10^{-4}$ at $z\sim 0.1$. Here $\mathcal{A}_{\rm H}$ should be understood as an order-of-magnitude estimate for the spatial component of the observer 4–acceleration, expressed in terms of the typical rate of change of peculiar velocities over a Hubble time.
This suggests that the required scale is not obviously negligible from a cosmological point of view, but rather comparable to cosmological acceleration scales associated with structure formation. The main discriminant is therefore not the amplitude alone, but the redshift evolution and directional stability of the dipole.
6 Observational Implementation and Testable Signatures
@@
$1+z\simeq(1+\bar{z})\left[1+(v_{s}-v_{o})_{\alpha}e^{\alpha}\right],$ (6.1)
@@
$\delta z\sim(1+\bar{z})\left[(v_{s}-v_{o})_{\alpha}e^{\alpha}+\int_{\lambda_{o}}^{\lambda_{s}}E\,\mathcal{A}_{\alpha}e^{\alpha}\,d\lambda\right],$ (6.2)
@@
$\delta z(\bar{z})=\,\hat{n}\cdot\mathbf{D}(\bar{z}),$ (6.3)
@@
$\mathbf{D}(\bar{z})=\mathbf{D}_{\rm kin}(\bar{z})+\mathbf{D}_{\rm ng}(\bar{z}),$ (6.4)
@@
Therefore, by measuring the redshift dependence of the dipole amplitude and its directional stability, one can test whether the observed signal is consistent with purely kinematic expectations or requires an additional non–geodesic contribution. This provides a concrete observational avenue to test whether the redshift dipole is fully accounted for by large–scale structure kinematics, or whether additional contributions are present.
7 Discussion and Conclusions
@@
In this work, it has been shown that non-geodesic observer congruences generically induce a dipolar modulation in the cosmological redshift through the projection of the four-acceleration along the line of sight. This effect arises at the level of the redshift itself and therefore propagates to any observable expressed in redshift space. The observable is defined operationally through redshift measurements in a specified frame (e.g. CMB frame), and therefore the effect corresponds to a physically distinct inference, rather than a coordinate artifact or a trivial reparametrization.
@@
A key aspect of this result is that the induced dipole is not equivalent to the standard kinematic dipole generated by a constant peculiar velocity. In the latter case, the dipole can be removed by a single global Lorentz boost. In contrast, the contribution derived here is sourced by the (generally position-dependent) acceleration of the observer congruence and can accumulate along the past lightcone, making it generically non-reducible to a global boost transformation and leading to a genuinely non-local contribution to the observed redshift.
@@
This point is closely related, but not equivalent, to previous work on tilted cosmologies, where observer congruences moving with respect to the matter flow exhibit modified kinematical properties (see e.g. [13, 1]). In such approaches, the focus is typically on local frame transformations and their impact on dynamical variables. By contrast, the effect derived here arises from the propagation of redshift along the past lightcone and therefore encodes an integrated contribution of the observer acceleration, which is not captured by local boosts alone.
This distinction has direct implications for cosmological inference. In practice, observational data are interpreted in terms of redshifts defined with respect to specific frames, such as the CMB frame. If such frames correspond to non-geodesic congruences, the inferred cosmological parameters may inherit systematic anisotropies or biases associated with observer kinematics. Standard treatments of redshift corrections in peculiar-velocity analyses (e.g. [4]) typically assume geodesic observer congruences, and therefore do not account for additional line-of-sight contributions associated with non-geodesic effects.
The order-of-magnitude estimates presented here indicate that the acceleration required to produce observable dipolar effects is comparable to typical cosmological acceleration scales associated with large-scale structure formation. This suggests that the effect may be relevant in low-redshift datasets, where peculiar velocities and bulk flows are significant. In this regime, the observable of interest is not only the dipole amplitude itself, but its redshift evolution and directional stability, which can in principle distinguish between kinematic and non-geodesic contributions. In particular, large-scale dipoles inferred from redshift surveys (e.g. [15]) provide a natural observational benchmark against which such effects could be tested. While these measurements are commonly interpreted in terms of bulk flows induced by large-scale structure, the framework presented here suggests that part of the observed signal could, in principle, receive contributions from the kinematics of the observer congruence itself.
@@
A quantitative assessment requires extending the present framework to realistic cosmological models and datasets, including forward modeling of redshift-space observables and their angular dependence. This can be implemented using peculiar-velocity catalogues (e.g., CosmicFlows) and, at higher redshift, forthcoming large surveys with wide sky coverage.
@@
More broadly, these results emphasize that cosmological observables are not determined solely by spacetime geometry, but also by the kinematics of the observer congruence used to interpret them. This highlights the importance of a fully covariant, observer-based approach to cosmology, in which the role of the observer is treated as an integral part of the physical description rather than a passive element.

### 14 Apr 2026 v2

@@ IV Non-geodesic observer congruence and the dipole modulation
$\frac{1}{1+z}\frac{d(1+z)}{d\lambda}=-EH-E\,\sigma_{\alpha\beta}e^{\alpha}e^{\beta}+E\,\mathcal{A}_{\alpha}e^{\alpha}.$ (36)
In the present work, we focus on the direction-dependent modulation of the redshift and therefore restrict attention to the anisotropic contributions. Under this assumption, and keeping the FLRW background relation for the scale factor evolution (30), the standard relation $(1+z)a=a_{0}$ is replaced by,
@@
$(1+z_{\rm obs})=\frac{a_{0}}{a}\exp\left[\int_{\lambda_{0}}^{\lambda_{s}}E\left(\mathcal{A}_{\alpha}e^{\alpha}-\sigma_{\alpha\beta}e^{\alpha}e^{\beta}\right)d\lambda\right].$ (37)
@@
For weak departures from the geodesic FLRW case, this becomes
@@
$1+z_{\rm obs}\simeq\frac{a_{0}}{a}\left[1+\int_{\lambda_{0}}^{\lambda_{s}}E\left(\mathcal{A}_{\alpha}e^{\alpha}-\sigma_{\alpha\beta}e^{\alpha}e^{\beta}\right)d\lambda\right].$ (38)
Since $\mathcal{A}_{\alpha}e^{\alpha}$ is the projection of a spatial vector along the line of sight, it has the angular structure of a dipole. Likewise, the contraction $\sigma_{\alpha\beta}e^{\alpha}e^{\beta}$ is the projection of a symmetric trace-free tensor and therefore has the angular structure of a quadrupole. This shows that non-geodesicity of the observer congruence naturally induces a dipolar modulation in the redshift itself, while anisotropic expansion encoded in the shear contributes a quadrupolar component. Therefore, the dipole and quadrupole anisotropies in redshift-space observables may provide complementary probes of the acceleration and shear of the observer congruence, respectively.
@@
The non-geodesic dipole arises from a line-of-sight integral of the observer acceleration field and therefore depends on the photon trajectory and redshift. This contrasts with the standard kinematic dipole, which is generated by a constant boost and is independent of distance. As a result, this dipole cannot, in general, be removed by a single global Lorentz transformation, since it arises from an integrated, position-dependent acceleration field rather than from a constant boost.
The shear term represents the anisotropic expansion of the chosen observer congruence. In an exact FLRW description, this term vanishes identically. However, in realistic inhomogeneous settings, an observer congruence used to interpret cosmological observables need not remain shear-free along the past light cone. In this sense, the quadrupolar contribution can be interpreted as encoding deviations from an idealized isotropic Hubble flow, potentially associated with the anisotropic kinematics of large-scale structure. This suggests a possible connection between low-redshift quadrupolar anisotropies in distance-redshift observables [14] and the anisotropic kinematics of nearby large-scale structure [6]. Exploring this observational connection in detail is beyond the scope of the present work and is left for future study.
V Upper Bound Estimate of the Non–Geodesic Dipole
The dipolar contribution associated with a non–geodesic observer congruence arises from the term $\mathcal{A}_{\alpha}e^{\alpha}$ in the covariant propagation equation 34. As discussed in the previous section, the different kinematical terms have distinct angular structures: the expansion scalar contributes to the monopole, the acceleration term $\mathcal{A}_{\alpha}e^{\alpha}$ generates a dipole, and the shear term $\sigma_{\alpha\beta}e^{\alpha}e^{\beta}$ induces a quadrupole. In this section, we focus on isolating the dipolar contribution associated with $\mathcal{A}_{\parallel}\equiv\mathcal{A}_{\alpha}e^{\alpha}$,
$\frac{d}{d\lambda}\ln(1+z)=-E\left(\frac{1}{3}\theta-\mathcal{A}_{\alpha}e^{\alpha}\right).$ (39)
To estimate the amplitude of the anisotropic modulation, we separate the monopole contribution and define
$\ln(1+z)\simeq\ln(1+\bar{z})+\delta\ln(1+z)_{\rm ng},$ (40)
where $\bar{z}$ is the monopolar redshift, satisfying
$\frac{d}{d\lambda}\ln(1+\bar{z})=-EH.$ (41)
The dipolar residual is then governed by
@@ V Upper Bound Estimate of the Non–Geodesic Dipole
$\left.\frac{d}{d\lambda}\,\delta\ln(1+z)\right|_{\rm ng}\simeq E\,\mathcal{A}_{\parallel}.$ (42)
@@
$k^{\alpha}=\frac{dx^{\alpha}}{d\lambda},$ (43)
and therefore
$E=-u_{\alpha}k^{\alpha}=\frac{dt}{d\lambda}.$ (44)
Using
$\frac{d}{d\lambda}=E\frac{d}{dt},$ (45)
Eq. (42) becomes
$\left.\frac{d}{dt}\delta\ln(1+z)\right|_{\rm ng}\simeq\mathcal{A}_{\parallel},$ (46)
so that, restoring dimensions,
$\left.\frac{d}{dt}\delta\ln(1+z)\right|_{\rm ng}\sim\frac{\mathcal{A}_{\parallel}}{c}.$ (47)
@@
$\delta\ln(1+z_{obs})_{\rm ng}\sim\int\frac{\mathcal{A}_{\parallel}}{c}\,dt.$ (48)
At this stage, it is important to emphasize that the evolution of the effective four-acceleration associated with an observer congruence in an inhomogeneous universe is currently a subject of ongoing debate. Different approaches, particularly relativistic and quasi–Newtonian treatments of large–scale structure formation, have led to qualitatively different conclusions regarding both the magnitude and time dependence of $\mathcal{A}_{a}$ (see e.g. [8, 2, 18]). In this work, we do not attempt to model the detailed evolution of $\mathcal{A}_{\parallel}(t)$. Instead, we adopt a conservative phenomenological approach aimed at estimating the possible amplitude of the induced dipole.
As a simplifying approximation, we consider the case where $\mathcal{A}_{\parallel}$ can be treated as approximately constant over a restricted low–redshift interval. This assumption over a restricted interval should be interpreted as an effective bound on the line-of-sight integral, rather than as a statement about the time evolution of the acceleration itself.
Under this approximation,
$\delta\ln(1+z_{obs})_{\rm ng}\lesssim\frac{\mathcal{A}_{\parallel}}{c}\,\Delta t\approx\frac{\mathcal{A}_{\parallel}L}{c^{2}},$ (49)
@@
$L\simeq\frac{c}{H_{0}}\,z.$ (50)
Substituting Eq. (50) into Eq. (49) yields
$\delta\ln(1+z_{obs})_{\rm ng}\lesssim\frac{\mathcal{A}_{\parallel}}{cH_{0}}\,z.$ (51)
For $z\ll 1$, one has $\delta\ln(1+z)\simeq\delta z$, so that
$\delta z_{\rm ng}\lesssim\frac{\mathcal{A}_{\parallel}}{cH_{0}}\,z.$ (52)
This relation can be recast as
$\mathcal{A}_{\parallel}\gtrsim\frac{cH_{0}}{z}\,\delta z_{\rm ng}.$ (53)
@@
$cH_{0}\simeq 6.8\times 10^{-10}\ {\rm m\,s^{-2}}.$ (54)
At $z\simeq 0.1$, representative of the shallow redshift regime relevant for peculiar-velocity surveys, this gives
$\mathcal{A}_{\parallel}\gtrsim 6.8\times 10^{-9}\,\delta z_{\rm ng}\quad{\rm m\,s^{-2}}.$ (55)
Therefore, dipole amplitudes in the range $\delta z_{\rm ng}\sim 10^{-3}$$10^{-5}$ correspond to accelerations
$\mathcal{A}_{\parallel}\gtrsim 10^{-12}-10^{-14}\ {\rm m\,s^{-2}}.$ (56)
For comparison, one may consider characteristic rates of change of peculiar velocities associated with large-scale structure formation, as discussed in both Newtonian and relativistic treatments (see e.g. [9]). These analyses suggest that the evolution of peculiar velocities occurs over cosmological timescales, leading to a dimensional scaling of the form
$\mathcal{A}_{\rm LSS}\sim vH_{0},$ (57)
where $v\sim 300\,{\rm km\,s^{-1}}$ is a typical peculiar velocity amplitude.
This scaling should be understood as a dimensional proxy for the rate of change of velocities over cosmological timescales, rather than as a prediction of a coherent acceleration field or as implying $\mathcal{A}_{\rm LSS}\sim vH_{0}$ at all times. In realistic structure formation scenarios, the evolution of peculiar velocities is driven by local gravitational potentials and is neither spatially uniform nor constant in time. This gives
$\mathcal{A}_{\rm LSS}\sim 7\times 10^{-13}\ {\rm m\,s^{-2}}.$ (58)
Using this as a reference scale, one finds that dipolar modulations of order $\delta z\lesssim 10^{-4}$ at $z\sim 0.1$ are compatible with this dimensional estimate, although a realistic prediction requires specifying the temporal evolution and coherence properties of the effective acceleration field.
The estimates presented here should be interpreted as order-of-magnitude upper bounds over a restricted low-redshift interval, aimed at assessing the potential significance of the effect rather than providing a detailed physical model.
VI Observational Implementation and Testable Signatures
@@
$1+z\simeq(1+\bar{z})\left[1+(v_{s}-v_{o})_{\alpha}e^{\alpha}\right],$ (59)
@@
$\delta z\sim(1+\bar{z})\left[(v_{s}-v_{o})_{\alpha}e^{\alpha}+\int_{\lambda_{o}}^{\lambda_{s}}E\,\mathcal{A}_{\alpha}e^{\alpha}\,d\lambda\right],$ (60)
@@
$\delta z(\bar{z})=\,\hat{n}\cdot\mathbf{D}(\bar{z}),$ (61)
@@
$\mathbf{D}(\bar{z})=\mathbf{D}_{\rm kin}(\bar{z})+\mathbf{D}_{\rm ng}(\bar{z}),$ (62)
@@
Therefore, by measuring the redshift dependence of the dipole amplitude and its directional stability, one can test whether the observed signal is consistent with purely kinematic expectations or requires an additional non–geodesic contribution. In practice, detecting this effect requires disentangling it from observational uncertainties and intrinsic scatter in peculiar velocity measurements, which may dominate at low redshift. Overall, this provides a concrete observational avenue to test whether the redshift dipole is fully accounted for by large–scale structure kinematics, or whether additional contributions are present.
VII Discussion and Conclusions
@@
In this work, we have shown that non-geodesic observer congruences generically induce a dipolar modulation in the cosmological redshift through the projection of the four-acceleration along the line of sight. This effect arises directly at the level of the redshift and therefore propagates to any observable expressed in redshift space. Since redshift is operationally defined with respect to a chosen observational frame (e.g. the CMB frame), this contribution corresponds to a physically distinct inference, rather than to a coordinate artifact or a trivial reparametrization.
@@
A key aspect of this result is that the induced dipole is not equivalent to the standard kinematic dipole generated by a constant peculiar velocity. In the latter case, the dipole can be removed by a single global Lorentz boost. In contrast, the contribution derived here is sourced by the generally position-dependent acceleration of the observer congruence and accumulates along the past lightcone. As a result, it cannot, in general, be removed by a global boost transformation and instead represents a genuinely non-local contribution to the observed redshift.
It is important to emphasize that this does not challenge the standard practice of applying local kinematic corrections to define observational frames such as the CMB rest frame. Rather, it highlights that such corrections do not uniquely define a geodesic observer congruence along the past lightcone, and therefore need not correspond to a physically global Hubble flow in an inhomogeneous universe. This highlights that redshift is not a purely geometric observable, but a kinematically defined quantity that depends on the choice of observer congruence. Consequently, cosmological inference based on redshift measurements implicitly inherits assumptions about the underlying kinematics of the chosen observational frame.
The covariant formulation also shows that different kinematical quantities generate distinct angular structures. While the acceleration term produces a dipolar modulation, the shear contribution induces a quadrupolar component. The latter could be related with anisotropic large-scale structure and may be related to low-redshift quadrupolar patterns in distance-redshift observables. A quantitative observational study of this connection is beyond the scope of the present work and will be explored in future analyses.
These considerations are particularly relevant in the low-redshift regime, where structure formation leads to significant inhomogeneities, peculiar velocity fields, and anisotropic environments. In this context, any effective non-geodesic contribution should be interpreted as a local and evolving property of the observer congruence. Accordingly, the estimates presented in this work are best regarded as order-of-magnitude upper bounds over a restricted low-redshift interval, rather than as a description of a coherent acceleration field or a detailed dynamical model.
@@
The distinction presented here is closely related, but not equivalent, to previous work on tilted cosmologies, where observer congruences moving with respect to the matter flow exhibit modified kinematical properties (see e.g. [17, 1]). In those approaches, the focus is typically on local frame transformations and their impact on dynamical variables. By contrast, the effect derived here arises from the propagation of redshift along the past lightcone and therefore encodes an integrated contribution of the observer kinematics that is not captured by local boosts alone.
From an observational perspective, this framework suggests that the relevant signature is not only the amplitude of a dipolar component, but also its redshift evolution and directional behaviour. A purely kinematic dipole associated with peculiar velocities is expected to trace the structure of the local velocity field, whereas a non-geodesic contribution may exhibit a distinct redshift dependence or a misalignment with respect to the reconstructed kinematic dipole. These features could, in principle, be tested using low-redshift peculiar-velocity catalogues (e.g. CosmicFlows) as well as distance-redshift data.
@@
The upper bound estimates indicate that the acceleration required to produce observable dipolar effects is comparable, at the level of dimensional analysis, to characteristic acceleration scales associated with large-scale structure formation. This suggests that the effect may be relevant in low-redshift datasets, where peculiar velocities and bulk flows are significant, although a quantitative prediction requires specifying the temporal evolution and coherence properties of the effective acceleration field. A quantitative assessment of these effects requires extending the present framework to realistic cosmological models and datasets, including forward modeling of redshift-space observables and their angular structure. This could be implemented using peculiar-velocity catalogues and, at higher redshift, forthcoming wide-field surveys.
It is interesting to note that the effect derived here is purely kinematical, yet it raises the question of how it relates to more familiar integrated contributions such as gravitational redshift effects in inhomogeneous spacetimes. While the present formulation does not rely on a potential-based description, clarifying the connection between these viewpoints could provide further insight into the physical origin of redshift anisotropies and deserves separate investigation.
@@
More broadly, these results highlight that cosmological observables are determined not only by spacetime geometry, but also by the kinematics of the observer congruence used to interpret them. This reinforces the importance of a fully covariant, observer-based approach to cosmology, in which the role of the observer is treated as an integral part of the physical description.